Laserfiche WebLink
Meeting Minutes <br /> Detailed, Grouped by Topic for each Meeting <br /> Item Meeting Item Description Resp Status Due Date Compl'd Cls'd <br /> 027-0 Reference 5/17/06-11/22/06 minutes for past Bob Heggenes No <br /> 02 discussions. (LCI) <br /> 12/7/2906: NBBJ sent BobH preliminary plans to <br /> provide ROM. <br /> 12/1312006: BobH to contact Blythe and Ewing for <br /> feedback,and provide price and write up tomorrow for <br /> OB unit changes. <br /> 12/20/2006: All AHU's will be on site by the end of <br /> January-Phase I & II. I.H.would like to review impact of <br /> Imaging changes prior to kicking off any OB changes. <br /> 1/3/07: OB decision to remain on hold until Lydig <br /> receives drawings and full impact of imaging can be <br /> reviewed. <br /> 2/14/2007: OB design could get started. Roof Drainage <br /> is on hold until OB is settled. Nurse call to be worked <br /> out in OB CP. <br /> Automatic Doors <br /> 031-0 1/10/07: Doors should be on emergency power at all Barbara Ringhouse No <br /> 08 times,and then set for either fail secure or fail safe for (ISLHOS);Jim <br /> medical and ambulance personnel. Brown(ISLHOS) <br /> 1/17/07: CarIT has confirmed function of all automatic <br /> doors with subcontractor and has e-mail verification out <br /> to confirm desired configuration. <br /> 1/24/07: CarIT returned submittal. NBBJ conversation <br /> with Fire Dept.on requirements of access. NBBJ may <br /> need to set up meeting with I.H.and Fire <br /> discuss usage. The 3 doors affected are the Main <br /> Lobby,ER entrance and Ambulance entrance. <br /> 1/31/07: BobH sent ROM to MarcE for review. <br /> interested in moving forward with this. NBBJ to review <br /> scope and issue CP. <br /> 2/7/2007: MarcE is putting together a meeting with Barb <br /> and Dan for further discussions. NBBJ to issue CP. <br /> ATS Alternate PCO <br /> 034-0 1/24/07: Lydig forwarded pricing for CP-30 to all. Rowe Moyle Yes <br /> 01 Sparling has reviewed and commented back to NBBJ (SPAELE) <br /> and I.H. This item does affect layout of electrical room <br /> in basement. Sparling has communicated to I.N.that <br /> footprint for the AB would be there. Lydig to review with <br /> Ewing. <br /> 1/31/07: Lydig has housekeeping pads. <br /> 2/7/2007: TomB had some discussions with Rowe <br /> yesterday,and will contact Mark and Vince for further <br /> review. BobH is waiting on a response from Rowe, <br /> since there has been some changes in equipment. <br /> NBBJ is working on CP-30. <br /> 2/14/2007: MarcE received CP and forwarded to Vince <br /> for signature. <br /> Prolog Manager Printed on: 2/15/2007 Lydig_Projects Page 2 <br />